OpenStudy (anonymous):

i need help with linear funcitions unit test algebra 1 foundations

OpenStudy (anonymous):

The answer is C

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

please i need the whole test i have to past it

OpenStudy (anonymous):

are you in connections academy?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It sucks doesn't it?

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

give me some questions.

OpenStudy (anonymous):

i give medals if you help

OpenStudy (anonymous):

so whats your questions. I'm in Algebra 2 so i've already done algebra 1

OpenStudy (anonymous):

what is the slope of the line that passes through the pair (2,9) and (5,12) -0 -1 -3 undefined

OpenStudy (anonymous):

algebra 1

OpenStudy (anonymous):

I believe it is either -1 or undefined, but let me check to be sure which one

OpenStudy (anonymous):

it is 1. because you calculate the change in the y's over the change in the x's, hence rise/run. however, when I do that, I get positive 1 not negative. This is a weird question.

OpenStudy (anonymous):

thanks i got alot more can you help

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what is the slope of the line that passes through the pair of points (-5.2,8.7) and (-3.2,2.7) -2 -2 - -2 - -3

OpenStudy (anonymous):

I got -3.

OpenStudy (anonymous):

write y=-0.4x+ 0.3 in standard form using intergers -4x+10y=3 =4x+10=-3 4x-10y=3 -4x+10y=3

OpenStudy (anonymous):

Actually for problems like these you can plug them into a website calculation tool called wolfram, and it will solve it for you. heres the website: http://www.wolframalpha.com/ Just type in the equation and say standard form next to it.

OpenStudy (anonymous):

But if wolfram doesn't work, my guess is B